# Can language models boost the power of randomized experiments without statistical bias?

### v1

@@ 2.2 CALM with few-shot learning
$\mathcal{S}^{*}(\mathcal{I}_{\ell_{1}}):=\big\{(X_{j,\ell_{1}}^{*},Z_{j,\ell_{1}}^{*},T^{*}_{j,\ell_{1}},Y_{j,\ell_{1}}^{*}),j=1,\ldots,m\big\},$
@@
(CALM estimator for $\mathbb{E}[Y(t)]$ with few-shot learning) We define for $i\in\mathcal{I}_{\ell_{3}}$:
@@
2.3 Estimation strategy of the heterogeneous calibration weight function
@@
Yet, in practice, the machine learning methods used to estimate the conditional means $\mu_{t}(x)$ and $\mu_{t}^{\dagger}(x)$ may rely on underlying modeling assumptions that deviate from the true data-generating mechanism, a phenomenon referred to as model misspecification [7, 60, 68]. Such discrepancies can arise in parametric settings, where the assumed functional form (e.g., linearity) is incorrect, as well as in nonparametric settings, where structural restrictions (such as piecewise linearity in tree-based approaches) are imposed but fail to capture the true conditional means [12, 25]. To address this potential model misspecification, we introduce a robust calibration strategy that accounts for heterogeneity in predictive power across subjects while guaranteeing efficiency gains relative to the classical AIPW estimator.
@@ 3 CALM for estimating other causal parameters
In this section, we extend the CALM framework introduced in Section 2 for estimating $\mathbb{E}[Y(t)]$ to accommodate other important causal parameters, including the average treatment effect (ATE) $\tau_{t,t^{\prime}}$ and the conditional average treatment effect (CATE) $\tau_{t,t^{\prime}}(x)$ between two treatment arms $t$ and $t^{\prime}$.
@@ 3.2 CALM for estimating CATE
Building on the ATE estimation in Section 3.1, we now describe estimation of the CATE $\tau_{t,t^{\prime}}(x)$. A standard AIPW-based CATE estimator regresses the influence function estimators $\widehat{\varphi}_{t,\texttt{AIPW}}(Y_{i},T_{i},X_{i})$ and $\widehat{\varphi}_{t^{\prime},\texttt{AIPW}}(Y_{i},T_{i},X_{i})$ on the covariates $X$ [31] (or on a subset of $X$; see 1). For concreteness, suppose the regression is implemented using a generic linear smoother. Then the AIPW-based CATE estimator can be written as
@@ 4.1 Theoretical properties of CALM with zero-shot learning
$\mathsf{V}_{t,\mathtt{CALM}}$ $=\mathbb{E}\left[\left(\mu_{t}(X)-\mu_{t}\right)^{2}+\frac{\mathbb{V}\mathrm{ar}(Y(t)\mid X)}{e_{t}(X)}\left(1-(1-e_{t}(X))\rho_{t}^{2}(X)\right)\right]$
$=\mathsf{V}_{t,\mathtt{AIPW}}-\mathbb{E}\left[\frac{\mathbb{V}\mathrm{ar}(Y(t)\mid X)}{e_{t}(X)}(1-e_{t}(X))\rho_{t}^{2}(X)\right]$
@@
Next, we examine to what extent the conclusions of Theorem 1 continue to hold without Assumption 3. Due to the Neyman orthogonality of our estimating equation, the CALM estimator remains consistent and asymptotically normally distributed. However, the asymptotic variance will take a different form. As expected, it is generally not possible to claim efficiency gains over the AIPW estimator when the nuisance functions used in our method are misspecified. Nevertheless, as we show in the remainder of this section, the CALM estimator remains (weakly) more efficient as long as the AIPW estimator uses the same misspecified models for the nuisance conditional expectations, and the heterogeneous calibration weights are specified in the robust form with a discrete representation of $X$ as described in Section 2.3. We now impose a high-level condition on potentially misspecified conditional mean estimators, analogous to those considered in [7] and [28].
@@
$\sigma_{t}^{2}(x^{\mathtt{C}}):=\mathbb{E}\big[\lambda_{t}(X)\{Y(t)-\widetilde{\mu}_{t}(X)\}^{2}\mid X^{\mathtt{C}}=x^{\mathtt{C}}\big],\quad\sigma_{t}^{\dagger 2}(x^{\mathtt{C}}):=\mathbb{E}\big[\lambda_{t}(X)\{Y^{\dagger}(t)-\widetilde{\mu}_{t}^{\dagger}(X)\}^{2}\mid X^{\mathtt{C}}=x^{\mathtt{C}}\big],$
@@
$\widetilde{\mathsf{V}}_{t,\mathtt{CALM}}$ $=\mathbb{E}\!\left[(\mu_{t}(X)-\mu_{t})^{2}+\mathbb{V}\mathrm{ar}(Y(t)\mid X)+\sigma_{t}^{2}(X^{\mathtt{C}})\{1-\widetilde{\rho}_{t}^{2}(X^{\mathtt{C}})\}\right]$
$=\widetilde{\mathsf{V}}_{t,\mathtt{AIPW}}-\mathbb{E}\!\left[\sigma_{t}^{2}(X^{\mathtt{C}})\widetilde{\rho}_{t}^{2}(X^{\mathtt{C}})\right],$
@@
$\widetilde{\rho}_{t}(x^{\mathtt{C}})=\frac{\mathbb{E}\!\left[\lambda_{t}(X)\{Y(t)-\widetilde{\mu}_{t}(X)\}\{Y^{\dagger}(t)-\widetilde{\mu}_{t}^{\dagger}(X)\}\,\middle|\,X^{\mathtt{C}}=x^{\mathtt{C}}\right]}{\sqrt{\sigma_{t}^{2}(x^{\mathtt{C}})\,\sigma_{t}^{\dagger 2}(x^{\mathtt{C}})}}.$
@@ 4.2 Theoretical properties of CALM with few-shot learning
$\bar{f}_{\theta,\mathtt{FS}}((x,z),t):=\mathbb{E}[Y^{\dagger}(t;\mathcal{S})\mid X=x,Z=z],$
where, as the notation suggests, the expectation is taken with respect to the random sample $\mathcal{S}$. The subscript $\theta$ represents the fixed parameters of the LLM. We then define the nuisance functions $\mu^{\dagger}_{t,\mathtt{FS}}(\cdot)$ and $\omega_{t,\mathtt{FS}}(\cdot)$ by replacing $Y^{\dagger}(t)$ in Step 3 with $\bar{f}_{\theta,\texttt{FS}}((X,Z),t)$; that is, $\mu^{\dagger}_{t,\mathtt{FS}}(x)=\mathbb{E}[\bar{f}_{\theta,\texttt{FS}}((X,Z),t)\mid X=x]$ and
@@
$\mathsf{V}_{t,{\mathtt{CALM},\mathtt{FS}}}$ $=\mathbb{E}\left[\left(\mu_{t}(X)-\mu_{t}\right)^{2}+\frac{\mathbb{V}\mathrm{ar}(Y(t)\mid X)}{e_{t}(X)}\left(1-(1-e_{t}(X))\rho_{t,\mathtt{FS}}^{2}(X)\right)\right]$
$=\mathsf{V}_{t,\mathtt{AIPW}}-\mathbb{E}\left[\frac{\mathbb{V}\mathrm{ar}(Y(t)\mid X)}{e_{t}(X)}(1-e_{t}(X))\rho_{t,\mathtt{FS}}^{2}(X)\right]$
@@ 4.3 Theoretical properties of CALM for other causal parameters
$\mathsf{V}_{t,t^{\prime},\mathtt{CALM}}$ $=\mathbb{E}\Big[(\tau_{t,t^{\prime}}(X)-\tau_{t,t^{\prime}})^{2}+\frac{\mathbb{V}\mathrm{ar}(Y(t)\mid X)}{e_{t}(X)}+\frac{\mathbb{V}\mathrm{ar}(Y(t^{\prime})\mid X)}{e_{t^{\prime}}(X)}$
$\qquad\;-\mathbb{C}\mathrm{ov}(V,Z\mid X)^{\prime}\Sigma_{V}^{-1}(X)\mathbb{C}\mathrm{ov}(V,Z\mid X)\Big]$
$=\mathsf{V}_{t,t^{\prime},\mathtt{AIPW}}-\mathbb{E}\!\left[\mathbb{C}\mathrm{ov}(V,Z\mid X)^{\prime}\Sigma_{V}^{-1}(X)\mathbb{C}\mathrm{ov}(V,Z\mid X)\right].$
@@
Suppose the assumptions of Corollary 2 hold. Further assume that the linear smoother weights $w_{i}(x;\mathbf{X})$ in Section 3.2 are generated from kernel weights $\frac{K\!\left(\frac{X_{i}-x}{h}\right)}{\sum_{j=1}^{n}K\!\left(\frac{X_{j}-x}{h}\right)}$ with bandwidth $h$ satisfying $h\to 0$, $nh^{p}\to\infty$, and $nh^{p+4}\to 0$ for covariate dimension $p$. Then, under standard conditions for kernel regression [36, 61, 66], for any fixed $x\in\mathcal{X}$, we have $\widehat{\tau}_{t,t^{\prime},\mathtt{CALM}}(x)\xrightarrow{p}\tau_{t,t^{\prime}}(x)$, and
@@ 5 Simulation studies
In this section, we evaluate the performance of CALM based on synthetic data calibrated from the BRIGHTEN study [45], which is an RCT with both structured and unstructured data. Before detailing our complete simulation design and results, we briefly summarize the key findings from our simulation studies as follows: First, CALM-based methods consistently exhibit higher estimation efficiency compared to AIPW-based methods, while maintaining statistical inferential validity (Figure 3). Second, when AIPW is augmented with few-shot-learning-based counterfactual predictions as covariates, it shows substantial bias and leads to coverage probabilities declining as the number of few-shot examples grows, leading to systematic under-coverage (Figures 3D). Third, the heterogeneous calibration weighting mechanism in CALM selectively leverages LLM-based predictions where they align most closely with observed outcomes. This adaptive borrowing mechanism across covariate strata results in substantial variance reduction compared to other benchmark methods.
@@ 5.1 Overview of BRIGHTEN study
The BRIGHTEN study provides a unique data structure that combines both structured covariates and rich unstructured textual baseline surveys (see Figure 1A for an overview), making it particularly suitable for evaluating CALM. To set the stage for our simulation design, which involves generating synthetic data resembling the BRIGHTEN study, we clarify key notations and their interpretation. We define a binary treatment variable, $T_{i}$, indicating assignment either to the intervention group receiving internet-based Problem-Solving Therapy (iPST), denoted as $T_{i}=1$, or to a control group assigned to alternative interventions, denoted as $T_{i}=0$. We include structured covariates $X_{i}$ such as demographic and clinical characteristics: sex, education, employment status, marital status, race, and age, etc. The unstructured textual covariates $Z_{i}$ capture participants’ self-reported reasons for enrollment and their satisfaction with the mobile app. The outcome of interest, $Y_{i}$, is the PHQ-9 score, a validated measure of depressive symptom severity with higher values indicating more severe depression symptoms [33].
@@ 5.2 Simulating BRIGTEN study with structured/unstructured data
Second, we learn both a generative model for the covariates and predictive models for the outcomes. We first model the joint covariate distribution $p(W)$ using a conditional Generative Adversarial Network (CTGAN) [69]. From the BRIGHTEN dataset, we construct a CTGAN training table that includes both the structured covariates and the two text-embedding blocks obtained from preprocessing. To improve sample efficiency and training stability, each 384-dimensional text block is compressed via principal component analysis (PCA) to 32 components.The structured covariates and PCA-compressed embeddings are then concatenated to form a unified training table, whose schema is validated with Metadata before fitting a CTGANSynthesizer [43]. This yields a generator $G$ that models $\widehat{p}(W)$, which we use to draw synthetic covariates for downstream simulations. Next, using the embedded BRIGHTEN dataset $\{W_{i},T_{i},Y_{i}\}_{i=1}^{n}$, we fit causal forests to learn the conditional mean and conditional variance, $\mu_{t}(W):=\mathbb{E}[Y|W,T=t]$, $\sigma_{t}^{2}(W):=\mathbb{E}[(Y-\mu_{t}(W)^{2})]$, $t\in\{0,1\}$. using the R package grf [58]. We denote the fitted models as $\widehat{\mu}_{t}(W)$ and $\widehat{\sigma}_{t}^{2}(W)$. Finally, because the original outcome is the PHQ-9 score (0-27), we categorize it into five clinically standard levels of depression severity—minimal, mild, moderate, high, and severe—to simplify prediction tasks for LLMs. LLMs are more effective when predicting discrete, semantically meaningful categories rather than continuous numeric scores. After classifying participants into these categories, we recode the categories numerically from 1 to 5 to represent increasing severity.

### v2

@@ 2.2 CALM with few-shot learning
$\mathcal{S}^{*}(\mathcal{I}_{\ell_{1}}):=\big\{(X_{j,\ell_{1}}^{*},Z_{j,\ell_{1}}^{*},T^{*}_{j,\ell_{1}},Y_{j,\ell_{1}}^{*}),j=1,\ldots,m\big\}$
@@
(CALM estimator for $\mathbb{E}(Y(t))$ with few-shot learning) We define for $i\in\mathcal{I}_{\ell_{3}}$:
@@
2.3 Estimation strategy of the heterogeneous calibration weights
@@
Yet, in practice, the machine learning methods used to estimate the conditional means $\mu_{t}(x)$ and $\mu_{t}^{\dagger}(x)$ may rely on underlying modeling assumptions that deviate from the true data-generating mechanism, a phenomenon referred to as model misspecification. Such discrepancies can arise in parametric settings, where the assumed functional form (e.g., linearity) is incorrect, as well as in semi-/non-parametric settings with additional structural restrictions (such as piecewise linearity or additive separability). To address this potential model misspecification, we introduce a robust calibration strategy that accounts for heterogeneity in predictive power across subjects while guaranteeing efficiency gains relative to the classical AIPW estimator.
@@ 3 CALM for estimating other causal parameters
In this section, we extend the CALM framework introduced in Section 2 for estimating $\mathbb{E}(Y(t))$ to accommodate other important causal parameters, including the average treatment effect (ATE) $\tau_{t,t^{\prime}}$ and the conditional average treatment effect (CATE) $\tau_{t,t^{\prime}}(x)$ between two treatment arms $t$ and $t^{\prime}$.
@@ 3.2 CALM for estimating CATE
Building on the ATE estimation in Section 3.1, we now describe estimation of the CATE $\tau_{t,t^{\prime}}(x)$. A standard AIPW-based CATE estimator regresses the influence function estimators $\widehat{\varphi}_{t,\texttt{AIPW}}(Y_{i},T_{i},X_{i})$ and $\widehat{\varphi}_{t^{\prime},\texttt{AIPW}}(Y_{i},T_{i},X_{i})$ on the covariates $X$ [26] (or on a subset of $X$; see 2). For concreteness, suppose the regression is implemented using a generic linear smoother. Then the AIPW-based CATE estimator can be written as
@@ 4.1 Theoretical properties of CALM with zero-shot learning
$\mathsf{V}_{t,\mathtt{CALM}}$ $=\mathbb{E}\left(\left(\mu_{t}(X)-\mu_{t}\right)^{2}+\frac{\mathbb{V}\mathrm{ar}(Y(t)\mid X)}{e_{t}(X)}\left(1-(1-e_{t}(X))\rho_{t}^{2}(X)\right)\right)$
$=\mathsf{V}_{t,\mathtt{AIPW}}-\mathbb{E}\left(\frac{\mathbb{V}\mathrm{ar}(Y(t)\mid X)}{e_{t}(X)}(1-e_{t}(X))\rho_{t}^{2}(X)\right)$
@@
Next, we examine to what extent the conclusions of Theorem 1 continue to hold without Assumption 3. Due to the Neyman orthogonality of our estimating equation, the CALM estimator remains consistent and asymptotically normally distributed. However, the asymptotic variance will take a different form. As expected, it is generally not possible to claim efficiency gains over the AIPW estimator when the nuisance functions used in our method are misspecified. Nevertheless, as we show in the remainder of this section, the CALM estimator remains (weakly) more efficient as long as the AIPW estimator uses the same misspecified models for the nuisance conditional expectations, and the heterogeneous calibration weights are specified in the robust form with a discrete representation of $X$ as described in Section 2.3. We now impose a high-level condition on potentially misspecified conditional mean estimators, analogous to those considered in [25].
@@
$\sigma_{t}^{2}(x^{\mathtt{C}}):=\mathbb{E}\big(\lambda_{t}(X)\{Y(t)-\widetilde{\mu}_{t}(X)\}^{2}\mid X^{\mathtt{C}}=x^{\mathtt{C}}\big),\quad\sigma_{t}^{\dagger 2}(x^{\mathtt{C}}):=\mathbb{E}\big(\lambda_{t}(X)\{Y^{\dagger}(t)-\widetilde{\mu}_{t}^{\dagger}(X)\}^{2}\mid X^{\mathtt{C}}=x^{\mathtt{C}}\big),$
@@
$\widetilde{\mathsf{V}}_{t,\mathtt{CALM}}$ $=\mathbb{E}\!\left((\mu_{t}(X)-\mu_{t})^{2}+\mathbb{V}\mathrm{ar}(Y(t)\mid X)+\sigma_{t}^{2}(X^{\mathtt{C}})\{1-\widetilde{\rho}_{t}^{2}(X^{\mathtt{C}})\}\right)$
$=\widetilde{\mathsf{V}}_{t,\mathtt{AIPW}}-\mathbb{E}\!\left(\sigma_{t}^{2}(X^{\mathtt{C}})\widetilde{\rho}_{t}^{2}(X^{\mathtt{C}})\right),$
@@
$\widetilde{\rho}_{t}(x^{\mathtt{C}})=\frac{\mathbb{E}\!\left(\lambda_{t}(X)\{Y(t)-\widetilde{\mu}_{t}(X)\}\{Y^{\dagger}(t)-\widetilde{\mu}_{t}^{\dagger}(X)\}\,\middle|\,X^{\mathtt{C}}=x^{\mathtt{C}}\right)}{\sqrt{\sigma_{t}^{2}(x^{\mathtt{C}})\,\sigma_{t}^{\dagger 2}(x^{\mathtt{C}})}}.$
@@ 4.2 Theoretical properties of CALM with few-shot learning
$\bar{f}_{\theta,\mathtt{FS}}((x,z),t):=\mathbb{E}(Y^{\dagger}(t;\mathcal{S})\mid X=x,Z=z),$
where, as the notation suggests, the expectation is taken with respect to the random sample $\mathcal{S}$. The subscript $\theta$ represents the fixed parameters of the LLM. We then define the nuisance functions $\mu^{\dagger}_{t,\mathtt{FS}}(\cdot)$ and $\omega_{t,\mathtt{FS}}(\cdot)$ by replacing $Y^{\dagger}(t)$ in Step 3 with $\bar{f}_{\theta,\texttt{FS}}((X,Z),t)$; that is, $\mu^{\dagger}_{t,\mathtt{FS}}(x)=\mathbb{E}(\bar{f}_{\theta,\texttt{FS}}((X,Z),t)\mid X=x)$ and
@@
$\mathsf{V}_{t,{\mathtt{CALM},\mathtt{FS}}}$ $=\mathbb{E}\left(\left(\mu_{t}(X)-\mu_{t}\right)^{2}+\frac{\mathbb{V}\mathrm{ar}(Y(t)\mid X)}{e_{t}(X)}\left(1-(1-e_{t}(X))\rho_{t,\mathtt{FS}}^{2}(X)\right)\right)$
$=\mathsf{V}_{t,\mathtt{AIPW}}-\mathbb{E}\left(\frac{\mathbb{V}\mathrm{ar}(Y(t)\mid X)}{e_{t}(X)}(1-e_{t}(X))\rho_{t,\mathtt{FS}}^{2}(X)\right)$
@@ 4.3 Theoretical properties of CALM for other causal parameters
$\mathsf{V}_{t,t^{\prime},\mathtt{CALM}}$ $=\mathbb{E}\Big((\tau_{t,t^{\prime}}(X)-\tau_{t,t^{\prime}})^{2}+\frac{\mathbb{V}\mathrm{ar}(Y(t)\mid X)}{e_{t}(X)}+\frac{\mathbb{V}\mathrm{ar}(Y(t^{\prime})\mid X)}{e_{t^{\prime}}(X)}$
$\qquad\qquad\;-\mathbb{C}\mathrm{ov}(V,Z\mid X)^{\prime}\Sigma_{V}^{-1}(X)\mathbb{C}\mathrm{ov}(V,Z\mid X)\Big)$
$=\mathsf{V}_{t,t^{\prime},\mathtt{AIPW}}-\mathbb{E}\!\left(\mathbb{C}\mathrm{ov}(V,Z\mid X)^{\prime}\Sigma_{V}^{-1}(X)\mathbb{C}\mathrm{ov}(V,Z\mid X)\right).$
@@
Suppose the assumptions of Corollary 2 hold. Further assume that the linear smoother weights $w_{i}(x;\mathbf{X})$ in Section 3.2 are generated from kernel weights $\frac{K\!\left(\frac{X_{i}-x}{h}\right)}{\sum_{j=1}^{n}K\!\left(\frac{X_{j}-x}{h}\right)}$ with bandwidth $h$ satisfying $h\to 0$, $nh^{p}\to\infty$, and $nh^{p+4}\to 0$ for covariate dimension $p$. Then, under standard conditions for kernel regression [31, 54], for any fixed $x\in\mathcal{X}$, we have $\widehat{\tau}_{t,t^{\prime},\mathtt{CALM}}(x)\xrightarrow{p}\tau_{t,t^{\prime}}(x)$, and
@@ 5 Simulation studies
In this section, we evaluate the performance of CALM based on synthetic data calibrated from the BRIGHTEN study [42], which is an RCT with both structured and unstructured data. Before detailing our complete simulation design and results, we briefly summarize the key findings from our simulation studies as follows: First, CALM-based methods consistently exhibit higher estimation efficiency compared to AIPW-based methods, while maintaining statistical inferential validity (Figure 3). Second, when AIPW is augmented with few-shot-learning-based counterfactual predictions as covariates, it shows substantial bias and leads to coverage probabilities declining as the number of few-shot examples grows, leading to systematic under-coverage (Figures 3D). Third, CALM’s efficiency gain is present regardless of whether unstructured covariates are incorporated, suggesting that the improvement does not arise simply from adding extra unstructured information but from intrinsic information gain encoded in the pre-trained LLMs (Figure 4). Lastly, the heterogeneous calibration weighting mechanism in CALM selectively leverages LLM-based predictions where they align most closely with observed outcomes. This adaptive borrowing mechanism across covariate strata results in substantial variance reduction compared to other benchmark methods (Figure 5).
@@ 5.1 Overview of BRIGHTEN study
The BRIGHTEN study provides a unique data structure that combines both structured covariates and rich unstructured textual baseline surveys, making it particularly suitable for evaluating CALM. Figure 1(A) summarizes the trial’s collected data, where we observe that unstructured data are collected at the initiation of the study. Figure 1(B) shows that baseline free-text responses on motivation for app use and follow-up app-satisfaction comments exhibit the highest cosine similarity with discretized PHQ-9 outcomes (trial primary outcome), surpassing demographics and app-usage metrics. Figure 1(B) demonstrates the potential of the unstructured data in the BRIGHTEN study as valuable “prognostic factors” (pre-treatment variables that are predictive of the outcome).
To set the stage for our simulation design, which involves generating synthetic data resembling the BRIGHTEN study, we clarify key notations and their interpretation. We define a binary treatment variable, $T_{i}$, indicating assignment either to the intervention group receiving internet-based Problem-Solving Therapy (iPST), denoted as $T_{i}=1$, or to a control group assigned to alternative interventions, denoted as $T_{i}=0$. We include structured covariates $X_{i}$ such as demographic and clinical characteristics: sex, education, employment status, marital status, race, and age, etc. The unstructured textual covariates $Z_{i}$ capture participants’ self-reported reasons for enrollment and their satisfaction with the mobile app. The outcome of interest, $Y_{i}$, is the PHQ-9 score, a validated measure of depressive symptom severity with higher values indicating more severe depression symptoms [28].
@@ 5.2 Simulating BRIGTEN study with structured/unstructured data
Second, we learn both a generative model for the covariates and predictive models for the outcomes. We first model the joint covariate distribution $p(W)$ using a conditional Generative Adversarial Network (CTGAN) [56]. From the BRIGHTEN dataset, we construct a CTGAN training table that includes both the structured covariates and the two text-embedding blocks obtained from preprocessing. To improve sample efficiency and training stability, each 384-dimensional text block is compressed via principal component analysis (PCA) to 32 components.The structured covariates and PCA-compressed embeddings are then concatenated to form a unified training table, whose schema is validated with Metadata before fitting a CTGANSynthesizer [40]. This yields a generator $G$ that models $\widehat{p}(W)$, which we use to draw synthetic covariates for downstream simulations. Next, using the embedded BRIGHTEN dataset $\{W_{i},T_{i},Y_{i}\}_{i=1}^{n}$, we fit causal forests to learn the conditional mean and conditional variance, $\mu_{t}(W):=\mathbb{E}(Y|W,T=t)$, $\sigma_{t}^{2}(W):=\mathbb{E}((Y-\mu_{t}(W)^{2}))$, $t\in\{0,1\}$. using the R package grf [50]. We denote the fitted models as $\widehat{\mu}_{t}(W)$ and $\widehat{\sigma}_{t}^{2}(W)$. Finally, because the original outcome is the PHQ-9 score (0-27), we categorize it into five clinically standard levels of depression severity—minimal, mild, moderate, high, and severe—to simplify prediction tasks for LLMs. LLMs are more effective when predicting discrete, semantically meaningful categories rather than continuous numeric scores. After classifying participants into these categories, we recode the categories numerically from 1 to 5 to represent increasing severity.